Circular No. 6514

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SATURN, URANUS AND NEPTUNE H. Feuchtgruber, Infrared Space Observatory (ISO) Science Opera-

tions Center and Max-Planck-Institut für Extraterrestrische Physik; T. de Graauw, Space Research Organization of the Netherlands and Kapteyn Astronomical Institute; E. Lellouch, B. Bezard and T. Encrenaz, Département de Recherche Spatiale, Observatoire de Paris; and M. J. Griffin, Queen Mary and Westfield College, report: “During October and November we observed Saturn, Uranus and Neptune in the 14–45-μm range with the ISO Short Wavelength Spectrometer in grating mode at a spectral resolution of 1000–1500. On Uranus, eight H_{2}O emission features were detected (28.43, 28.59, 33.00, 35.91, 39.38, 40.34, 43.89 and 44.19 μm); a very preliminary analysis suggests that these lines are formed at temperatures of 200–300 K in the microbar region of the planet’s upper atmosphere. Six H_{2}O lines were also detected on Neptune, and three were found on Saturn. In addition, the Q branch of the CO_{2} ν_{2} band at 14.98 μm was observed in emission in the spectra of Saturn and Neptune.”

GRO J2058+42 C. A. Wilson, NASA Marshall Space Flight Center; T. Strohmayer,

NASA Goddard Space Flight Center and Universities Space Research Association; and D. Chakrabarty, Massachusetts Institute of Technology, report: “The Compton Observatory/BATSE instrument has observed a series of five outbursts from the 195.6-s accreting x-ray pulsar GRO J2058+42, with a bright initial outburst in Sept. 1995 (IAUC 6238) followed by four weaker (15–20-mCrab pulsed flux in the 20–50-keV band) outbursts, all spaced by about 110 days. The weak outbursts have each lasted about two weeks, with a marginal decline in peak intensity over consecutive outbursts. The most recent outburst peaked on 1996 Nov. 25 ± 2 with a 20–50-keV pulsed flux of 14 ± 1 mCrab. The barycentric pulse frequency on Nov. 25.0 UT was 5.1125±0.0001 mHz. A target-of-opportunity scan of the Compton/OSSE/BATSE error box (IAUC 6239) on Nov. 28 with the Proportional Counter Array (PCA) on the Rossi X-ray Timing Explorer (RXTE) constrained the 90-percent-confidence region for the source position to a circle with 4′ radius and centered at α = 20^{h}59^{m}.0, δ = +41^{o}43′ (equinox 2000.0). RXTE/PCA observed a peak flux of 11 ± 1 mCrab (2–30 keV) and a pulse frequency consistent with the BATSE measurement. Optical observations of the new error circle are strongly encouraged in order to identify the optical companion, most likely a Be star.”

1996 December 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden